OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 8, -16, 32, -64,...

OpenStudy (anonymous):

A. an = 8(-2)n-1 B. an = 2(-8)n-1 C. an = -8(2)n-1

OpenStudy (anonymous):

\[a _{n} = a _{1}r ^{n - 1}\]

OpenStudy (anonymous):

are you sure? can you explain how you got to that conclusion?

OpenStudy (rogue):

As Calcmathlete said, the formula for a geometric sequence is given by\[a_n = a_1 r^{n-1}\]where a1 is your first term and r is your common ratio between each consecutive term. In the sequence you have, your first term is 8, so that is your value for a1. If you divide the 2nd term by the 1st term, or the 3rd by the 2nd, or any term by the term before it, you get -2, which is your common ratio. So the formula for your sequence is\[a_n = 8(-2)^{n-1}\]Choice A.

OpenStudy (anonymous):

Thank you so much. That is so helpful. I couldn't understand it for the life of me!